# MedXChat: Bridging CXR Modalities with a Unified Multimodal Large Model

### v1

@@ 3.3 Multi Model Tasks
The proposed framework is a unified model shown in Fig. 1 bottom right, integrating three distinct tasks: CXR-to-Report, CXR-VQA and Text-to-CXR, elaborated below.
CXR-to-Report   In this task, a pre-trained CLIP visual encoder ViT-L/14 [41] is used to process the CXR image, providing text-aligned visual features. These visual features are then paired with a corresponding user’s instruction, e.g., ”Generate a report of this chest x-ray”, and input into the LLM. Based on this instruction, LLM produces a concise report according to the information extracted from the CXR.
CXR-VQA   In this task, our model inputs a set of three conditions into the LLM: the CXR image, the accompanying medical report, and a question from the user like ”Is there cardiomegaly?”. The LLM evaluates these inputs and responds with an answer, for instance, ”No, cardiomegaly is not present.”
@@
”Human: Hi there, I recently had a chest X-ray and the report mentioned ”right lower lobe consolidation”. I’m not sure what that means. Could you generate a lateral view of a chest X-ray with that condition for me?”
@@ 3.4 Training Objectives
Our training objective encompasses two components: LLM at the instruction tuning stage and CXR-SD at the preparation stage. LLM’s objective is to produce comprehensive target paragraphs consisting of an explanation, input, and an autoregressive response, akin to the instruction-following fine-tuning paradigm [46] used in GPT pre-training [40, 5]. The loss function is specifically calculated based on the tokens generated in response to the prompt, aligning with the key response mechanism of the model. Moreover, the training objective of CXR-SD follows that of the SD model, focusing on denoising [43]: it employs textual prompts and temporal training steps to forecast the noise that has been introduced to the distorted images. Through this process, the model incrementally diminishes the noise, thereby progressively restoring the clarity of the image with each step.
Large Language Model. With the dialogues $\left(\mathbf{X}_{\mathrm{q}}^{1},\mathbf{X}_{\mathrm{a}}^{1},\cdots,\mathbf{X}_{\mathrm{q}}^{T},\mathbf{X}_{\mathrm{a}}^{T}\right)$ obtained utilizing ChatGPT-4 in Sec. 3.1, where $\mathbf{X}_{\mathrm{q}}$ is the question, $\mathbf{X}_{\mathrm{a}}$ is the answer, and T is the total number of dialogues, we finetune LLM with LoRA [22] through minimizing the loss $L_{a}$:
@@ 4.1 Dataset
MIMIC-CXR [25] stands as the largest publicly available dataset encompassing chest radiographs along with free-text reports. It covers 14 categories related to lung disease diagnosis. Part of the CXR images are obtained in both PA and lateral views. In our research, we followed the official dataset division guidelines of MIMIC-CXR to maintain consistency and enable fair comparisons. Consequently, our training dataset comprises 270,790 samples of image-report pairs, accompanied by sets of 2,130 and 3,858 samples for validation and testing, respectively.
Table 1: CXR-to-Report: AUROC and F1. †  indicates the results are quoted from [30].
@@
Table 2: CXR-to-Report: NLP Metrics. †  indicates the results are quoted from their respective papers.
@@ 4.3 Evaluation Metrics
CXR-VQA. We follow the VQA performance evaluation framework from ELIXR [58] to enable the comparison with other LLM-based methods [^1] This choice is necessitated by the unavailability of CXR-VQA results for other LLM-based methods. Specifically, from the MIMIC-CXR dataset, we randomly choose eight samples from each of the six distinct diagnostic categories. We either inquiry about specific lesions or query the presence, location, and severity of the findings in each CXR image.
Table 3: CXR-VQA: Accuracy by topic. †  indicates the results are quoted from [30]. “SST” stands for Size, Severity, Type.
@@
Table 4: CXR-VQA: Accuracy by Diagnosis. †  indicates the results are quoted from [30].
@@ 4.4 Results
CXR-to-Report. The results of AUROC and F1 are compared in Table 1, which are critical metrics to reflect the clinic efficacy of the generated reports. The results of the comparing methods are directly quoted from the paper of LLM-CXR [30]. As seen, our model produces the highest overall scores in terms of both F1 and AUROC among the compared methods. It is also the best/second performer across all diagnostic categories except ”No Fundings”. As also observed, the non-LLM-based UniXGen [29] underperforms when compared with LLM-based methods XrayGPT [47] and LLM-CXR [30], showing the potential of LLMs in generating clinic-relevant contents.
NLP metrics are given in Table 2. In addition to large multi-modal generation methods, we also include conventional methods exclusively for text generation. As seen, our model’s BLEU scores (BLEU-1 to BLEU-4) surpass nearly all methods, indicating a substantial overlap of n-grams in our generation and the ground truth. Moreover, compared with LLM-CXR, the sole method capable of performing all three tasks within a single integrated framework like ours, our model exhibits indisputable superiority across all NLP metrics except CIDEr. However, the notably high CIDEr of LLM-CXR might be linked to a distinct evaluation protocol focusing on the impression section rather than the complete report  [^2] CXR-to-Report performance evaluation of LLM-CXR using complete reports yields lower results than those reported in Table 2. Ours also outperforms UniXGen [29] in all BLEU scores and CIDEr. Notably, our model excels in generating clinically consistent reports as already seen in Table 1. Compared to conventional methods, ours leads in BLEU-1 to BLEU-3, with only a slight lag behind GSK in BLEU-4. Our performance is average in ROUGE, yet superior to a majority of models in METEOR and CIDEr.

### v2

@@ 3.3 Multi Model Tasks
The proposed framework is a unified framework shown in Fig. 1 bottom right, integrating three distinct tasks: CXR-to-Report, CXR-VQA and Text-to-CXR, elaborated below.
CXR-to-Report   In this task, a pre-trained CLIP visual encoder ViT-L/14 [40] is used to process the CXR image, providing text-aligned visual features. These visual features are then paired with a corresponding user’s instruction, e.g., "Generate a report of this chest x-ray", and input into the LLM. Based on this instruction, LLM produces a concise report according to the information extracted from the CXR.
CXR-VQA   In this task, our model inputs a set of three conditions into the LLM: the CXR image, the accompanying medical report, and a question from the user like "Is there cardiomegaly?". The LLM evaluates these inputs and responds with an answer, for instance, "No, cardiomegaly is not present."
@@
"Human: Hi there, I recently had a chest X-ray and the report mentioned "right lower lobe consolidation". I’m not sure what that means. Could you generate a lateral view of a chest X-ray with that condition for me?"
@@ 3.4 Training Objectives
Our training objective encompasses two components: LLM at the instruction tuning stage and CXR-SD at the preparation stage. LLM’s objective is to produce comprehensive target paragraphs consisting of an explanation, input, and an autoregressive response, akin to the instruction-following fine-tuning paradi-gm [46] used in ChatGPT pre-training [41, 4]. The loss function is specifically calculated based on the tokens generated in response to the prompt, aligning with the key response mechanism of the model. Moreover, the training objective of CXR-SD follows that of the SD model, focusing on denoising [43]: it employs textual prompts and temporal training steps to forecast the noise that has been introduced to the distorted images. Through this process, the model incrementally diminishes the noise, thereby progressively restoring the clarity of the image with each step.
Large Language Model. With the dialogues $\left(\mathbf{X}_{\mathrm{q}}^{1},\mathbf{X}_{\mathrm{a}}^{1},\cdots,\mathbf{X}_{\mathrm{q}}^{T},\mathbf{X}_{\mathrm{a}}^{T}\right)$ obtained utilizing ChatGPT-4 in Sec. 3.1, where $\mathbf{X}_{\mathrm{q}}$ is the question, $\mathbf{X}_{\mathrm{a}}$ is the answer, and T is the total number of dialogues, we fine-tune LLM with LoRA [22] through minimizing the loss $L_{a}$:
@@ 4.1 Dataset
MIMIC-CXR [25] stands as the largest publicly available dataset encompassing chest radiographs along with free-text reports. It covers 14 categories related to lung disease diagnosis. Part of the CXR images are obtained in both PA and lateral views. In our research, we followed the official dataset division guidelines of MIMIC-CXR to maintain consistency and enable fair comparisons. Consequently, our training dataset comprises 270,790 samples of image-report pairs, accompanied by 2,130 and 3,858 samples for validation and testing, respectively.
Table 1: CXR-to-Report: AUROC and F1. † marks quoted results from [30].
@@
Table 2: CXR-to-Report: NLP Metrics. † marks quoted results from respective papers.
@@ 4.3 Evaluation Metrics
CXR-VQA. We follow the VQA performance evaluation framework from ELIX-R [59] to enable the comparison with other LLM-based methods [^1] This choice is necessitated by the unavailability of CXR-VQA results for other LLM-based methods. Specifically, from the MIMIC-CXR dataset, we randomly choose eight samples from each of the six distinct diagnostic categories. We either inquiry about specific lesions or query the presence, location, and severity of the findings in each CXR image.
Table 3: CXR-VQA: Accuracy by topic. † marks quoted results from [30]. “SST" stands for Size, Severity, Type.
@@
Table 4: CXR-VQA: Accuracy by Diagnosis. † marks quoted results from [30].
@@ 4.4 Results
CXR-to-Report. The results of AUROC and F1 are compared in Table 1, which are critical metrics to reflect the clinic efficacy of the generated reports. The results of the comparing methods are directly quoted from the paper of LLM-CXR [30]. As seen, our model produces the highest overall scores in terms of both F1 and AUROC among the compared methods. It is also the best/second performer across all diagnostic categories except “No Fundings". As also observed, the non-LLM-based UniXGen [29] underperforms when compared with LLM-based methods XrayGPT [47] and LLM-CXR [30], showing the potential of LLMs in generating clinic-relevant contents.
NLP metrics are given in Table 2. In addition to large multimodal generation methods, we also include conventional methods exclusively for text generation. As seen, our model’s BLEU scores (BLEU-1 to BLEU-4) surpass nearly all methods, indicating a substantial overlap of n-grams in our generation and the ground truth. Moreover, compared with LLM-CXR, the sole method capable of performing all three tasks within a single integrated framework like ours, our model exhibits indisputable superiority across all NLP metrics except CIDEr. However, the notably high CIDEr of LLM-CXR might be linked to a distinct evaluation protocol focusing on the impression section rather than the complete report  [^2] CXR-to-Report performance evaluation of LLM-CXR using complete reports yields lower results than those reported in Table 2. Ours also outperforms UniXGen [29] in all BLEU scores and CIDEr. Notably, our model excels in generating clinically consistent reports as already seen in Table 1. Compared to conventional methods, ours leads in BLEU-1 to BLEU-3, with only a slight lag behind GSK in BLEU-4. Our performance is average in ROUGE, yet superior to a majority of models in METEOR and CIDEr.
@@
Table 6: Radiologist evaluation on 40 randomly selected generation cases
#prefer/total Reports CXRs
UniXGen [29] 1/20 7/20
LLM-CXR [30] 0/20 0/20
MedXChat (Ours) 19/20 13/20
@@
Radiologist Evaluation. To further validate the performance of the three tasks, we facilitated a detailed comparative evaluation by a radiologist, focusing on the CXR-to-Report and Text-to-CXR tasks, as exemplified in Fig.6. Acknowledging the impracticality of exhaustive manual evaluation due to the MIMIC-CXR dataset’s extensive size, we randomly selected a representative subset of 40 generative cases for comparison. This subset included results generated by us, UniXGen [29], and LLM-CXR [30], containing 20 reports and 20 frontal CXRs. The CXRs only featured the frontal view because UniXGen and LLM-CXR are capable of generating only frontal CXRs. The evaluations predominantly favored the reports and CXRs generated by our model, with only a minority preferring the outputs from UniXGen. This trend underscores our approach’s superior performance and its consistent preference in comparative analyses.